HUBBLE SPACE TELESCOPE DAILY REPORT #2375 PERIOD COVERED: 0000Z (UTC) 05/18/99 - 0000Z (UTC) 05/19/99 Daily Status Report as of 139/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7949 (CCD Dark and Bias Monitor -- Cont. to June= 99) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a CCD dark calibration and a bias monitoring program. It was also scheduled to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 7282 (A Continuing Snapshot Survey for Compact Stellar Groups and Clusters around Wolf-Rayet Stars) The WF/PC-2 was used to make observations of the Wolf-Rayet star WR145 as part of a continuing snapshot survey of compact stellar groups and clusters. The newly imaged clusters allow us to accurately determine the distances, luminosities, and ages of the WR members. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed STIS/CCD/MA1 7511 (Spectroscopy of the Interstellar Medium in the Magellanic Bridge) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make spectrographic observations of the interstellar medium (Section DGIK975) in the Magellanic Bridge. The proposed observations will provide an opportunity to investigate the physical conditions within the Bridge gas, and to determine the chemical composition of important elements such as C, N, O, Mg, Si, S, and Fe. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Two Sets of WF/PC-2 8441 (WF/PC-2 Decontaminations and Associated Observations Pt. 1/2) The WF/PC-2 was used to make its monthly decontamination operations. Also included are instrument monitors tied to decons: photometric stability check, focus monitor, pre- and post-decon internals {bias, internal flats, k-spots, & darks}, and ultraviolet throughput checks. All observations were executed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed WF/PC-2 6833 (A Cepheid Search in the Virgo Cluster Galaxy NGC 4571) The WF/PC-2 was used to make observations of the Virgo Cluster galaxy NGC-4571 in order to search for Cepheid variable stars. The determination of the distance to the Virgo Cluster is key to establishing the value of the H_zero. The observations were completed as planned, and no problems were reported. 1.8 Completed Two Sets of STIS/CCD 8470 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation V) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations as part of a POMS proposal. This is a STIS non-scripted parallel Condition VI proposal. As described in HSTAR 7134 and 2.1, the acquisition for the second iteration of this proposal defaulted to fine lock one one FGS only. Otherwise, the observations were completed as scheduled, and no further anomalies were reported. 1.9 Completed WF/PC-2 6461 (High-Resolution Next Bright Galactic Nova) The WF/PC-2 was used to make high resolution images of the bright galactic nova U-Scorpii after it erupted to search for fluorescence and reflection light echoes from, and small-scale structure in, the nova circumstellar gas and dust ejecta. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Two Sets of STIS/CCD 8406 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation V) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations as part of a POMS proposal. This is a STIS non-scripted parallel Condition V proposal. The observations were completed as scheduled, and no anomalies were reported. 1.11 Completed WF/PC-2 6604 (Photometry of the Oldest Field Population in the Magellanic Clouds.) The WF/PC-2 was used to make both visual and inferred photometric observations of selected stars in the oldest field in the Large Magellanic Cloud LMC-063600 (a.k.a., NGC-121) From these data, the ages and population densities of the cluster will be measured. As described in HSTAR 7134 and 2.1, the acquisition for this proposal defaulted to fine lock one one FGS only. Otherwise, the observations were completed as scheduled, and no further problems were noted. 1.12 Completed Two Sets of STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. As described in HSTAR 7134 and 2.1, the acquisition for this proposal defaulted to fine lock one one FGS only. Otherwise, the observations were completed as scheduled, and no further problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 As documented in HSTAR 7134, the acquisition at 139/170549Z defaulted to fine lock back-up on FGS-3 only when FGS-1 exceeded the scan step limit. The subsequent re-acquisitons at 138/1839Z, 138/1933Z, 138/2022Z, 138/2109, and 138/2205Z also defaulted to fine lock on one FGS only. The proposals described in 1.8, 1.11, and 1.12 may have been affected. Scheduled Reacquisitions : 10 Successful : 9 The acquisition at 139/012508Z failed when the scan step limit was exceeded on FGS-3. HSTAR 7135 was written. Two proposals, occurring during the period of the next report (STIS 8042 and STIS 7949) were affected. 2.2 FHST Updates: Scheduled : 17 Successful : 17 2.3 Operations Notes: The HST operations team remains in the SAMS back-up facility using the PRS ground system. PRS-2 was re-booted at 139/0107Z due to tape drive problems. Using ROP SR-1, the SSR EDAC error counter was cleared once. The =97=CD=CDC-1 status buffer was dumped at 138/1258Z per ROP NS-3. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH